OpenStudy (anonymous):

What is the simplified form of the quantity 9 x squared minus 25 over the quantity 3 x minus 5 ?

OpenStudy (anonymous):

\[\frac{ 9*(-25)^2 }{ 3*-5 }=\frac{ 9*-5 }{ -3 }=15\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

;)